IN a legend the hero kicked a baby pig so that he is projected with a speed greater than that of his cry. if the weight of the baby pig is assumed to be 5kg and the time of contact 0.01 sec ,then what is the force with which the hero kicked him?i think it is related as J = Favg t ..and J = p ..then p= Favg t ..and (mv - 0)= Favg t ...is the cry of the baby is related to the velocity of sound?


Grade:10

2 Answers

Arnav Preet
7 years ago
As,F=MA So, here weight is given then we have to take its mass by W=MG So M=0.5 A=v-u/tv=340 (as given speed of his cry or sound)u=0 (as initially at rest) a=34000m/s2 Now, F=0.5×34000=17000N
Michael
3 years ago
The baby pig is kicked so that its velocity is greater than that of his cry.
It means Force required to set the body in required motion will be greater than the force needed to set the baby in velocity v=330m/s
Now, m=5kg,v=330,u=0m/s,∆t=0.01s
Let required force be F
 
As per question
F>m(v-u)/∆t
putting the given value
F>1.65×10^5N
only one option (b) 2×10^5 conciliate this
ANSWER {B} CORRECT